6 cards: One A,Two B's,Three C's / Die:6 Sides

think about first picking a card and
then rolling the die. Give the probability in simplest form.
P(A, 1) = _____

Answer by stanbon(75887)

6 cards: One A,Two B's,Three C's / Die:6 Sides
think about first picking a card and
then rolling the die. Give the probability in simplest form.
P(A, 1) = _____
===========================
P(A and 1) = P(A)*P(1) because the events are independent
---
= (1/6)*(1/6) = 1/36
=============================
Cheers,
Stan H.